40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7716     Entry time: Thu Nov 15 21:52:48 2012
Author: Den, Ayaka 
Type: Update 
Category: Green Locking 
Subject: yarm locked 

We aligned and locked Y arm for green:

  • installed camera on PSL to monitor green transmission
  • aligned green path on the ETMY table to see the beam on the PSL camera
  • misaligned ETMY and aligned ITMY to see reflected beam on REFL PD
  • installed green transmission PD on PSL
  • aligned ETMY and locked YARM to 00 mode

I've switched error channel cable to output monitor. Whitening filter is need for scattering measurements.

   etmyf.png  yarm.png

ELOG V3.1.3-